Given that Ka for HCN is 6.2*10^-10, calculate the pH of a 0.15 M KCN solution.
Answer #1

Answer : Complete derivation of expression of pH of salt of strong base (KOH) and weak acid (HCN) is given in the image attached

pH = 11.19

For future use direct formula

pH = 1/2 (pKw + pKa + log C) = 1/2 ( 14 + 9.20 -0.82 ) = 11.19